# Towards Robust Model-Based Reinforcement Learning Against Adversarial Corruption

### v2

@@ 4.1 Algorithm: CR-OMLE
We propose Algorithm 1 for the online episodic learning setting. In each round $t$, after observing an initial state $x_{t}^{1}$, the agent follows the principle of optimism and selects the model $M_{t}$ from the confidence set $\mathcal{M}_{t}$ to maximize the value function $V_{M}^{1}(x_{t}^{1})$. Subsequently, the agent follows the greedy policy $\pi_{t}$ to collect a trajectory ${(x_{t}^{h},a_{t}^{h},r_{t}^{h})}{h=1}^{H}$. In constructing the confidence set, the agent initially learns $\bar{M}{t+1}$ by maximizing the weighted log-likelihood, where the weight $\sigma_{t}^{h}$ is a truncated variant of the Information Ratio (IR) (1), referred to as uncertainty
@@
$U_{t}(x_{t}^{h},a_{t}^{h})=\sup_{M\in\mathcal{M}_{t}}\frac{{\mathrm{TV}}\big{(}P_{M}^{h}(\cdot|x_{t}^{h},a_{t}^{h})\|P_{\bar{M}_{t}}^{h}(\cdot|x_{t}^{h},a_{t}^{h})\big{)}}{\sqrt{\lambda+\sum_{s=1}^{t-1}{\mathrm{TV}}\big{(}P_{M}^{h}(\cdot|x_{s}^{h},a_{s}^{h})\|P_{\bar{M}_{t}}^{h}(\cdot|x_{s}^{h},a_{s}^{h})\big{)}^{2}/\sigma_{s}^{h}}}.$ (4)
@@ 4.2 Regret Bounds
$\mathrm{Reg}(T)=$ $\tilde{\mathcal{O}}\Big{(}H\log B\sqrt{T\log|\mathcal{M}|{\mathrm{ED}}(\mathcal{M},\sqrt{\lambda/T})}+CH\cdot{\mathrm{ED}}(\mathcal{M},\sqrt{\lambda/T})\Big{)}.$
This regret bound consists of two parts. The main part $H(T\log B\log|\mathcal{M}|{\mathrm{ED}}(\mathcal{M},\sqrt{\lambda/T}))^{1/2}$ is unrelated to corruption and matches the bound for OMLE (Liu et al., 2023a, ). According to Theorem D.2, in the linear MDP setting with $d$ dimensions, ${\mathrm{ED}}(\mathcal{M},\sqrt{\lambda/T})=O(d)$. Thus, the regret reduces to $\tilde{\mathcal{O}}(H\log B\sqrt{Td\log|\mathcal{M}|}+CHd)$, where the corruption part $CHd$ also matches the lower bound from Theorem 4.2. We highlight the proof sketch in the sequel and delay the detailed proof to Appendix A.1.
@@
Since it is hard to know the corruption level ahead of time, we discuss the solution for the unknown corruption level. Note that only the choice of parameter $\alpha=\sqrt{\log|\mathcal{M}|\log^{2}B}/C$ requires the knowledge of $C$, so we following He et al., (2022) to replace $C$ in $\alpha$ by a predefined corruption tolerance threshold $\bar{C}$. As long as the actually corruption level $C$ is no more then the tolerance threshold $\bar{C}$, we can still obtain a non-trivial regret bound.
@@
Under the same conditions as Theorem 4.1, let $\beta_{t}^{h}=\Theta(\sqrt{\log(|\mathcal{M}|/\delta)\log^{2}B})$ and $\alpha=\sqrt{\log|\mathcal{M}|}/\bar{\zeta}$. If $C\leq\bar{C}$, we have with probability at least $1-\delta$,
$\mathrm{Reg}(T)=$ $\tilde{\mathcal{O}}\Big{(}H\log B\sqrt{T\log|\mathcal{M}|{\mathrm{ED}}(\mathcal{M},\sqrt{\lambda/T})}+\bar{C}H\cdot{\mathrm{ED}}(\mathcal{M},\sqrt{\lambda/T})\Big{)}.$
@@
[…]
@@
We emphasize key aspects of the proof in the following discussion and postpone the detailed explanation to Appendix C.1. The key point in in this step is to multiply the corruption term by a sufficiently small coefficient $\alpha$ to nullify the impact of the corruption $C$. It is important to note that the analysis for regression and Maximum Likelihood Estimation (MLE) differs significantly. While we can naturally separate the square error into corruption-unrelated and corruption-related terms and use uncertainty weights to control the latter in regression problems (He et al.,, 2022; Ye et al., 2023a, ; Ye et al., 2023b, ), in the MLE formulation, it is challenging not only to isolate the corruption terms but also to ensure their multiplication by corruption $c_{s}^{h}$ and uncertainty.
To begin with, we highlight the susceptibility of estimation and regret to the corruption level $C$ without incorporating uncertainty weights. The expected difference in log-likelihood between $M$ and $M_{*}$ can be expressed as:
$\sum_{s=1}^{t-1}{\mathbb{E}}_{x^{h+1}\sim P^{h}_{s}(\cdot|z_{s}^{h})}\log\sqrt{\frac{\mathrm{d}P^{h}_{\bar{M}_{t}}(x^{h+1}|z_{s}^{h})}{\mathrm{d}P^{h}_{*}(x^{h+1}|z_{s}^{h})}}$
$\qquad=\underbrace{\sum_{s=1}^{t-1}{\mathbb{E}}_{x^{h+1}\sim P^{h}_{*}(\cdot|z_{s}^{h})}\log\sqrt{\frac{\mathrm{d}P^{h}_{\bar{M}_{t}}(x^{h+1}|z_{s}^{h})}{\mathrm{d}P^{h}_{*}(x^{h+1}|z_{s}^{h})}}}_{\text{TV~{}Term}}+\underbrace{\sum_{s=1}^{t-1}{\mathbb{E}}_{x^{h+1}\sim(P^{h}_{s}-P^{h}_{*})(\cdot|z_{s}^{h})}\log\Big{(}\frac{\mathrm{d}P^{h}_{\bar{M}_{t}}(x^{h+1}|z_{s}^{h})}{\mathrm{d}P^{h}_{*}(x^{h+1}|z_{s}^{h})}\Big{)}}_{\text{Corruption~{}Term}}.$
From the definition of Hellinger distance and TV-norm, the TV term can be bounded by
$-\sum_{s=1}^{t-1}{\mathrm{TV}}(P^{h}_{*}(\cdot|z_{s}^{h})\|P^{h}_{\bar{M}_{t}}(\cdot|z_{s}^{h}))^{2}.$
Since the TV-norm is at most $1$, the corruption term is only bounded by
$\sum_{s=1}^{t-1}c_{s}^{h}{\mathrm{TV}}(P^{h}_{*}(\cdot|z_{s}^{h})\big{\|}P^{h}_{\bar{M}_{t}}(\cdot|z_{s}^{h}))\leq C.$
This leads to $\beta=\Theta(\log B\sqrt{\log(|\mathcal{M}|/\delta)}+C)$ and a $\tilde{\mathcal{O}}(C\sqrt{T})$ regret bound. This bound becomes trivial when $C=\Omega(\sqrt{T})$.
To deal with this problem, we bring in the uncertainty weights to obtain
${\mathbb{E}}\sum_{s=1}^{t-1}\frac{1}{\sigma_{s}^{h}}\log\sqrt{\frac{\mathrm{d}P^{h}_{\bar{M}_{t}}(x^{h+1}|z_{s}^{h})}{\mathrm{d}P^{h}_{*}(x^{h+1}|z_{s}^{h})}}\lesssim-\sum_{s=1}^{t-1}\frac{{\mathrm{TV}}\big{(}P^{h}_{*}(\cdot|z_{s}^{h})\big{\|}P^{h}_{\bar{M}_{t}}(\cdot|z_{s}^{h})\big{)}^{2}}{\sigma_{s}^{h}}+\underbrace{\sum_{s=1}^{t-1}\frac{c_{s}^{h}U_{s}(z_{s}^{h})\beta}{\sigma_{s}^{h}}}_{\text{Corruption~{}Term}},$
where we use $\lesssim$ to omit constants for conciseness, and $U_{s}$ is the uncertainty defined in (4). The corruption term can be further bounded by $\alpha C\sqrt{\lambda+\beta^{2}}$ since $\sigma_{s}^{h}\geq U_{s}(z_{s}^{h})/\alpha$. By similarly bounding the variance and applying the Bernstein inequality, we demonstrate that $M_{*}\in\mathcal{M}_{t}$ due to the non-negativity of the Total Variation (TV)-norm. Additionally, the in-sample error is upper-bounded by $\beta^{2}$ since $\bar{M}_{t}$ is the Maximum Likelihood Estimation (MLE).
@@
$\tilde{\mathcal{O}}\Bigg{(}\sqrt{TH\log|\mathcal{M}|\sum_{h=1}^{H}\sup_{\mathcal{S}_{T}}\sum_{t=1}^{T}{\mathbb{E}}^{t}_{\pi_{t}}\big{(}I^{h}_{\sigma}(\lambda,\mathcal{M}_{t},\mathcal{S}_{t})\big{)}^{2}}+C\sum_{h=1}^{H}\sup_{\mathcal{S}_{T}}\sum_{t=1}^{T}{\mathbb{E}}^{t}_{\pi_{t}}\big{(}I^{h}_{\sigma}(\lambda,\mathcal{M}_{t},\mathcal{S}_{t})\big{)}^{2}\Bigg{)},$
@@
$I^{h}_{\sigma}(\lambda,\mathcal{M},\mathcal{S}_{t}^{h})=\min\bigg{\{}1,\sup_{M\in\mathcal{M}_{t}}\frac{{\mathrm{TV}}(P_{M}^{h}(\cdot|z_{t}^{h})\|P_{\bar{M}_{t}}^{h}(\cdot|z_{t}^{h}))/(\sigma_{t}^{h})^{1/2}}{\sqrt{\lambda+\sum_{s=1}^{t-1}{\mathrm{TV}}(P_{M}^{h}(\cdot|z_{s}^{h})\|P_{\bar{M}_{t}}^{h}(\cdot|z_{s}^{h}))^{2}/\sigma_{s}^{h}}}\bigg{\}}.$ (5)
Moreover, to establish an instance-independent bound, we follow Ye et al., 2023a to demonstrate the relationship between the sum of weighted Information Ratio (IR) concerning dataset $\mathcal{S}_{t}={z_{s}^{h}}_{s=1}^{t}$ and the eluder dimension:
@@ 5.1 Algorithm: CR-PMLE
In the offline learning setting, we introduce our algorithm, Corruption-Robust Pessimistic MLE, shown in Algorithm 2. In this algorithm, we also estimate the model by maximizing the log-likelihood with uncertainty weights and constructing the confidence set $\widehat{\mathcal{M}}$. To address the challenge that weights cannot be computed iteratively in rounds, as in the online setting, we adopt the weight iteration algorithm (Algorithm 3) from Ye et al., 2023b to obtain an approximated uncertainty. They prove that this iteration will converge since the weights are monotonically increasing and upper-bounded. Moreover, as long as the weights $\sigma_{t}^{h}$ are in the same order as the truncated uncertainty, the learning remains robust to corruption. For completeness, we present the convergence result in Lemma B.1.
@@
$\sigma_{t}^{k}\leftarrow\max\Big{(}1,\sup_{M,M^{\prime}\in\mathcal{M}}\frac{l_{t}^{h}(M,M^{\prime})/\alpha}{\sqrt{\lambda+\sum_{s=1}^{T}l_{s}^{h}(M,M^{\prime})^{2}/\sigma_{s}^{k-1}}}\Big{)};$
@@
which shares a similar spirit with the model-free literature (Xie et al.,, 2021).

### v3

@@ 4 Online Model-based RL
4.1 Uncertainty Weighting for Model-based RL
In this subsection, we will illustrate how our uncertainty-weighting technique differs from the one for model-free RL. The main difficulty arises from the different estimators applied by these two settings. In model-free RL, we estimate the value functions using a value target regression. In contrast, for model-based RL, we directly estimate the hidden transition probability with MLE.
In detail, for model-free RL, let $f^{*}$ be the uncorrupted true value function and $\hat{f}$ be the least squares estimator with weights $\sigma_{s}$. The corruption term can be directly decomposed as the multiplication between uncertainty $U_{s}=\sup_{f\in\mathcal{F}}\frac{|\hat{f}(z_{s})-f(z_{s})|}{\sqrt{\lambda+\sum_{j=1}^{t}(\hat{f}(z_{j})-f(z_{j}))^{2}/\sigma_{j}}}$ and corruption (Ye et al., 2023a, b):
$\sum_{s=1}^{t}\frac{(\hat{f}(z_{s})-f^{*}(z_{s}))^{2}}{\sigma_{s}}=\sum_{s=1}^{t}\frac{(\hat{f}(z_{s})-y_{s})^{2}-(f^{*}(z_{s})-y_{s})^{2}}{\sigma_{s}}$
$\qquad+2\sum_{s=1}^{t}\frac{(\hat{f}(z_{s})-f^{*}(z_{s}))\epsilon_{s}}{\sigma_{s}}+2\underbrace{\sum_{s=1}^{t}\frac{U_{s}c_{s}}{\sigma_{s}}\cdot\beta}_{\text{Corruption~{} term}}.$
By choosing $\sigma_{s}\geq U_{s}/\alpha$, we can convert the corruption term to $\alpha C\beta$. To conclude, only if we transform the corruption term into the multiplication between uncertainty $U_{s}$ and corruption $c_{s}$, can the uncertainty-related weights cancel $U_{s}$ and bring in a small hyper-parameter $\alpha$.
While for model-based RL, it is difficult to decompose the corruption term as the multiplication between uncertainty and corruption terms from the log-likelihood, especially when uncertainty is based on the total variation (TV) norm. Specifically, let $P_{M}$ be the estimated probability, $P_{*}$ be the true probability, and $P_{t}$ be the corrupted probability at round $t$. Neglecting the upscript $h$ for convenience, we write:
$\mathbb{E}\frac{1}{\sigma_{s}}\log\sqrt{\frac{\mathrm{d}P_{\widehat{M}}(x|z_{s})}{\mathrm{d}P_{*}(x|z_{s})}}=\underbrace{\frac{1}{\sigma_{s}}\int\mathrm{d}P_{*}(x|z_{s})\log\sqrt{\frac{\mathrm{d}P_{\widehat{M}}(x|z_{s})}{\mathrm{d}P_{*}(x|z_{s})}}}_{\text{Uncorrupted~{}term}}$
$\qquad+\underbrace{\frac{1}{\sigma_{s}}\int(\mathrm{d}P_{s}(x|z_{s})-\mathrm{d}P_{*}(x|z_{s}))\log\sqrt{\frac{\mathrm{d}P_{\widehat{M}}(x|z_{s})}{\mathrm{d}P_{*}(x|z_{s})}}}_{\text{Corrupted~{}term}}.$
To make the corrupted term a multiplication between uncertainty and corruption, we use $\log x\leq x-1$ and decompose the integration into two regions according to whether $\mathrm{d}P_{s}-\mathrm{d}P_{*}$ is positive or not. Then, we deal with the variance similarly and use Assumption 3.1 and the Freedman inequality to obtain
${\mathbb{E}}\sum_{s=1}^{t-1}\frac{1}{\sigma_{s}^{h}}\log\sqrt{\frac{\mathrm{d}P^{h}_{\bar{M}_{t}}(x^{h+1}|z_{s}^{h})}{\mathrm{d}P^{h}_{*}(x^{h+1}|z_{s}^{h})}}$
$\lesssim-\sum_{s=1}^{t-1}\frac{{\mathrm{TV}}\big{(}P^{h}_{*}(\cdot|z_{s}^{h})\big{\|}P^{h}_{\bar{M}_{t}}(\cdot|z_{s}^{h})\big{)}^{2}}{\sigma_{s}^{h}}+\underbrace{\sum_{s=1}^{t-1}\frac{c_{s}^{h}U_{s}(z_{s}^{h})\beta}{\sigma_{s}^{h}}}_{\text{Corruption~{}Term}},$
where we use $\lesssim$ to omit constants for conciseness, and $U_{s}$ is the uncertainty defined in (4.2). Therefore, our new technique makes it possible for an TV-based uncertainty weight to control the corruption term.
@@ 4.2 Algorithm: CR-OMLE
We propose Algorithm 1 for the online episodic learning setting. In each round $t$, after observing an initial state $x_{t}^{1}$, the agent follows the principle of optimism and selects the model $M_{t}$ from the confidence set $\mathcal{M}_{t}$ to maximize the value function $V_{M}^{1}(x_{t}^{1})$. Subsequently, the agent follows the greedy policy $\pi_{t}$ to collect a trajectory $\{(x_{t}^{h},a_{t}^{h},r_{t}^{h})\}_{h=1}^{H}$. In constructing the confidence set, the agent initially learns $\bar{M}_{t+1}$ by maximizing the weighted log-likelihood, where the weight $\sigma_{t}^{h}$ is a truncated variant of the Information Ratio (IR) (1), referred to as uncertainty
@@
$U_{t}(x_{t}^{h},a_{t}^{h})$
$=\sup_{M\in\mathcal{M}_{t}}\frac{{\mathrm{TV}}\big{(}P_{M}^{h}(\cdot|x_{t}^{h},a_{t}^{h})\|P_{\bar{M}_{t}}^{h}(\cdot|x_{t}^{h},a_{t}^{h})\big{)}}{\sqrt{\lambda+\sum_{s=1}^{t-1}{\mathrm{TV}}\big{(}P_{M}^{h}(\cdot|x_{s}^{h},a_{s}^{h})\|P_{\bar{M}_{t}}^{h}(\cdot|x_{s}^{h},a_{s}^{h})\big{)}^{2}/\sigma_{s}^{h}}}.$ (4)
@@
Computation Efficiency. If the model class is finite or has a finite cover with cardinality $M$, the computational complexity is $MTH$ since in Line 4 and 7 of Algorithm 1, we have to search over also the components in the model class. If the model class is complicated, we have to acknowledge that methods like version space methods (Liu et al., 2023a; Jiang et al., 2017; Wang et al., 2023; Jin et al., 2021), which construct a confidence set, face computational drawbacks as they are computationally inefficient. However, practical algorithms may only need to leverage the insight that optimism is helpful and may not require such thorough exploration. Instead, we could construct a bonus and add it to the value function, and then choose the greedy policy for the optimistic value function (Curi et al., 2020).
When computing the weights, calculating the uncertainty quantity in practical scenarios is the main difficulty. Inspired by Ye et al. (2023b), we can approximate the uncertainty by ”bootstrapped uncertainty”. The intuition is that if the transition probability can be embedded into a vector space $P_{M}^{h}(x^{h+1}|z^{h})=\nu^{h}(M,x^{h+1})^{\top}\phi^{h}(z^{h})$. According to Example D.3, the uncertainty can be reduced to $\|\phi^{h}(z^{h})\|_{(\Sigma_{t}^{h})^{-1}}$, where the covariance matrix $\Sigma_{t}^{h}=\sum_{s=1}^{t-1}\phi^{h}(z_{s}^{h})\phi^{h}(z_{s}^{h})^{\top}$. Hence, from Ye et al. (2023b), the bootstrapped variance (called bootstrapped uncertainty) is an estimation of the uncertainty. To compute the bootstrapped uncertainty, we first learn $K$ transition probabilities independently with different seeds. Then, we take the uncertainty estimation as $\sqrt{\mathrm{Var}_{i=1,\ldots,K}[P_{M_{k}}(z)]}$. We leave the development of practical algorithms as future work.
@@ 4.3 Regret Bounds
$\quad+CH\cdot{\mathrm{ED}}(\mathcal{M},\sqrt{\lambda/T})\Big{)}.$
This regret bound consists of two parts. The main part $H(T\log B\log|\mathcal{M}|{\mathrm{ED}}(\mathcal{M},\sqrt{\lambda/T}))^{1/2}$ is unrelated to corruption and matches the bound for OMLE (Liu et al., 2023a). According to Theorem D.2, in the linear MDP setting with $d$ dimensions, ${\mathrm{ED}}(\mathcal{M},\sqrt{\lambda/T})=O(d)$. Thus, the regret reduces to $\tilde{\mathcal{O}}(H\log B\sqrt{Td\log|\mathcal{M}|}+CHd)$, where the corruption part $CHd$ also matches the lower bound from Theorem 4.2. We can achieve sub-linear regret when the corruption level is sub-linear. We highlight the proof sketch in the sequel and delay the detailed proof to Appendix A.1.
@@
Since it is hard to know the corruption level ahead of time, we discuss the solution for the unknown corruption level. Note that only the choice of parameter $\alpha=\sqrt{\log|\mathcal{M}|\log^{2}B}/C$ requires the knowledge of $C$, so we following He et al. (2022) to replace $C$ in $\alpha$ by a predefined corruption tolerance threshold $\bar{C}$. As long as the actually corruption level $C$ is no more then than the tolerance threshold $\bar{C}$, we can still obtain a non-trivial regret bound.
@@
Under the same conditions as Theorem 4.1, let $\beta_{t}^{h}=\Theta(\sqrt{\log(|\mathcal{M}|/\delta)\log^{2}B})$ and $\alpha=\sqrt{\log|\mathcal{M}|}/\bar{C}$. If $C\leq\bar{C}$, we have with probability at least $1-\delta$,
$\quad+\bar{C}H\cdot{\mathrm{ED}}(\mathcal{M},\sqrt{\lambda/T})\Big{)}.$
@@
To ensure that the true, uncorrupted model $M^{*}$ belongs to the confidence set $\mathcal{M}_{t}$, we demonstrate in the following lemma that $M^{*}$ satisfies (8). Moreover, for any $M\in\mathcal{M}_{t}$, the in-sample error is bounded.
@@
The key ideas of the analysis are presented in Subsection 4.1, and we postpone the detailed explanation to Appendix C.1.
@@
$\tilde{\mathcal{O}}\Bigg{(}$ $\sqrt{TH\log|\mathcal{M}|\sum_{h=1}^{H}\sup_{\mathcal{S}_{T}}\sum_{t=1}^{T}{\mathbb{E}}^{t}_{\pi_{t}}\big{(}I^{h}_{\sigma}(\lambda,\mathcal{M}_{t},\mathcal{S}_{t})\big{)}^{2}}$
$+C\sum_{h=1}^{H}\sup_{\mathcal{S}_{T}}\sum_{t=1}^{T}{\mathbb{E}}^{t}_{\pi_{t}}\big{(}I^{h}_{\sigma}(\lambda,\mathcal{M}_{t},\mathcal{S}_{t})\big{)}^{2}\Bigg{)},$
@@
$I^{h}_{\sigma}(\lambda,\mathcal{M},\mathcal{S}_{t}^{h})=\min\bigg{\{}1,\sup_{M\in\mathcal{M}_{t}}$
$\quad\frac{{\mathrm{TV}}(P_{M}^{h}(\cdot|z_{t}^{h})\|P_{\bar{M}_{t}}^{h}(\cdot|z_{t}^{h}))/(\sigma_{t}^{h})^{1/2}}{\sqrt{\lambda+\sum_{s=1}^{t-1}{\mathrm{TV}}(P_{M}^{h}(\cdot|z_{s}^{h})\|P_{\bar{M}_{t}}^{h}(\cdot|z_{s}^{h}))^{2}/\sigma_{s}^{h}}}\bigg{\}}.$ (5)
Moreover, to establish an instance-independent bound, we follow Ye et al. (2023a) to demonstrate the relationship between the sum of weighted Information Ratio (IR) concerning dataset $\mathcal{S}_{t}={z_{s}^{h}}_{s=1}^{t}$ and the eluder dimension:
@@ 5.1 Algorithm: CR-PMLE
In the offline learning setting, we introduce our algorithm, Corruption-Robust Pessimistic MLE, shown in Algorithm 2. In this algorithm, we also estimate the model by maximizing the log-likelihood with uncertainty weights and constructing the confidence set $\widehat{\mathcal{M}}$. To address the challenge that weights cannot be computed iteratively in rounds, as in the online setting, we adopt the weight iteration algorithm (Algorithm 3) from Ye et al. (2023b) to obtain an approximated uncertainty. They prove that this iteration will converge since the weights are monotonically increasing and upper-bounded. Moreover, as long as the weights $\sigma_{t}^{h}$ are in the same order as the truncated uncertainty, the learning remains robust to corruption. For completeness, we present the convergence result in Lemma B.1.
@@
$\sigma_{t}^{k}\leftarrow\max\Big{(}$ $1,\sup_{M,M^{\prime}\in\mathcal{M}}$
$\frac{l_{t}^{h}(M,M^{\prime})/\alpha}{\sqrt{\lambda+\sum_{s=1}^{T}l_{s}^{h}(M,M^{\prime})^{2}/\sigma_{s}^{k-1}}}\Big{)};$
@@
which shares a similar spirit with the model-free literature (Xie et al., 2021).